OpenStudy (anonymous):

Find the 3rd term of an arithmetic sequence with t2=9/2 and t5=6 A.) 3 B.) 5 C.) 2 D.) 4 I chose 3 and I got it wrong but I can't find where I made my mistake.

OpenStudy (anonymous):

If you could type it using the equation box it would be helpful.. I can't tell which ones are subscripts and which ones are actually part of the equation.

OpenStudy (anonymous):

Okie :) \[t _{2}=9/2\]

OpenStudy (anonymous):

and \[t _{5}=6\]

OpenStudy (anonymous):

I redid the problem is it 5?

OpenStudy (anonymous):

I'm working on it.

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

I think it would be 5 too, 9/2= 4.5 if you add .75 to it then t3= 5.25 and if you add .75 again then it would equal 6. I'm not quite sure if it's the problem but 5.25 is close to 5.

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

\[t_n=t_x+(n-x)*r\]\[t_5=t_2+(5-2)*r\]\[t_5-t_2=3r\]\[r=\frac{3}{2}\]\[t_n=t_2+(n-2)*\frac{3}{2}\]\[t_3=\frac{9}{2}+(3-2)*\frac{3}{2}\]\[t_3=\frac{9}{2}+\frac{3}{2}\]\[\boxed{t_3=6}\]

OpenStudy (anonymous):

Hold on, I made an mistake

OpenStudy (anonymous):

\[r=0.5\]\[t_n=t_2+(n-2)*0.5\]\[t_3=t_2+(3-2)*0.5\]\[t_3=\frac{9}{2}+0.5\]\[\boxed{\boxed{t_3=5}}\]